Circular No. 6043

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994W IN NGC 4041 C. Pollas, Observatoire de la Côte d’Azur, reports the following accu-

rate position for SN 1994W from a 5-min exposure taken on July 30.90 UT (cf. IAUC 6042): α = 11^{h}59^{m}37^{s}.67, δ = +62^{o}25′14′′.6 (equinox 1950.0); the corresponding offset from the center of the galaxy is 7′′.8 west and 17′′.5 north. The estimated magnitude of SN 1994W is m_{pv} ∼ 14. A nearby star (m_{pv} ∼ 14.7) has position end figures 53^{s}.68, 26′42′′.8.

M. W. Richmond, Princeton University; and R. R. Treffers, A. V. Filippenko, and S. D. Van Dyk, University of California at Berkeley, report that four R-band images of NGC 4041 were obtained in July as part of the Leuschner Observatory Supernova Search (observing conditions generally poor), yielding magnitudes with respect to a comparison star ∼ 94′′ east and 96′′ north of the nucleus (in the sense SN 1994W − comparison star): July 7, [+3.6; 13, [+3.1; 18, +1.2; 26, 0.0:. SN 1994W is measured as 6′′.7 west and 16′′.1 north of the galaxy’s nucleus.

P83l-57 A. W. Rodgers and W. H. Roberts, Mount Stromlo and Siding Spring

Observatories, report the discovery of a subdwarf O-type star with Balmerline and Ca II H- and K-line emission: “The emission-line systems have variable intensity and radial velocity. The star is in the Rodgers et al. catalogue of high-latitude, early-type stars (1993, A.J. 106, 591) and is designated there as P83l-57 (α = 3^{h}34^{m}34^{s}.23, δ = −64^{o}00′56′′.1, equinox 2000.0; magnitude V = 14.2). A series of slit spectra (range 385–438 nm, and sometimes to 510 nm; resolution 0.12 nm) have been obtained in 1991, 1992, and 1993. No well-defined periodicity in the emission-line intensity or velocity has been identified, but variations occur on a timescale of hours, and we encourage other observers to help obtain more complete coverage of these variations. The emission-line system presumably arises from an accretion disk associated with a companion in orbit about the subdwarf O star. The variations in the intensities of the Balmer emission do not correlate with those of the Ca II emission. The subdwarf O absorption spectrum appears constant, with lines of Fe III and He II detected, and an equivalent width of Hδ of 1.10 ± 0.07 nm, implying an effective temperature > 21 000 K. The heliocentric radial velocity is +160 ± 28 km/s. The emission-line system of the Balmer and Ca II H and K lines varies in radial velocity, with an amplitude of 70 km/s. The emission-line widths vary from unresolvable (< 60 km/s) to ∼ 160 km/s.”

1994 August 1 Daniel W. E. Green